# Faster Linear Systems and Matrix Norm Approximation via Multi-level Sketched Preconditioning

### v1

@@ 4 Two-Level MSP for Positive Definite Linear Systems
and for the other side we have
@@
For the second level we use preconditioned Lanczos to solve the $s\times s$ linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{C}^{\top}\mathbf{r}$. As for the preconditioner we choose $\mathbf{M}_{2}=\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\tilde{\lambda}\mathbf{W}=\mathbf{C}^{\top}\mathbf{\Phi}^{\top}\mathbf{\Phi}\mathbf{C}+\tilde{\lambda}\mathbf{W}$, where $\mathbf{\Phi}\in\mathbb{R}^{\phi\times n}$ is a sparse sketching matrix constructed using standard subspace embedding techniques [CDDR24] (see Lemma 4.4 below) with choice $\phi=O(s+\log 1/\delta)$. By Lemma 4.4, with probability $21-\delta/$ 2 we have $\kappa_{\mathbf{M}_{2}}\coloneqq\kappa(\mathbf{M}_{2}^{-1/2}(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{M}_{2}^{-1/2})=O(1)$. Since in Algorithm 1 we pre-compute $\mathbf{M}_{2}^{-1}$ exactly, the error caused by the preconditioner is 00. By using Theorem 6.1 we know that after $O(\sqrt{\kappa_{\mathbf{M}_{2}}}\log(\kappa_{\mathbf{M}_{2}}/\epsilon_{1}))=O(\log 1/\epsilon_{1})=O(\log(\kappa/\epsilon_{0}))$ iterations, preconditioned Lanczos will output $\hat{\mathbf{y}}$ that satisfies
@@ 4.2 Coarse Nyström Preconditioner based on Sparse Embedding
In this section we give a proof of Lemma 4.2, which measures the quality of Nyström preconditioner based on a sparse embedding matrix. Notice that compare with the most commonly used guarantee that the preconditioned linear system has a constant condition number, we are using a “coarse” Nyström preconditioner which benefits us by reducing the cost of constructing it, while slightly sacrificing its quality and allowing the condition number after preconditioning to be “not so small”.
The Nyström preconditioner we use is based on a kind of sketching matrices called sparse subspace embedding matrices. Recent work [CDDR24] gave the optimal dimension result for sparse oblivious subspace embedding (OSE), showing an optimal $O(d)$ result for embedding dimension.
Lemma 4.4 (Adapted from Theorem 1.4 in [CDDR24]).
However, the way [CDDR24] constructed the sparse embedding matrix $\mathbf{\Phi}$ does not apply to our case of Nyström preconditioner. Instead, following [CNW16, Coh16], we consider a more generally adopted sparse embedding matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$ defined as follows: the columns of $\mathbf{S}$ are independent, and for each column of $\mathbf{S}$, there are $\gamma$ random entries chosen uniformly without replacement and set to ±$\pm 1/\sqrt{\gamma}$ independently, with other entries in that column being set to 00. Notice that given $\mathbf{A}\in\mathbb{R}^{n\times n}$, $\mathbf{S}\mathbf{A}$ can be computed in time $O(\gamma\cdot\mathrm{nnz}(\mathbf{A}))$. To characterize the properties of sparse embedding matrix, we first introduce the following notion of oblivious subspace embedding moments property, which can be seen as a generalization of the “JL-moment property” introduced by [KN14].
@@
[Coh16] showed that sparse embedding matrix with $s=O(d\log(d/\delta)/\epsilon^{2})$ and $\gamma=O(\log(d/\delta)/\epsilon)$ satisfies $(\epsilon,\delta,d,\log(d/\delta))$-OSE moments property, as stated in the following lemma.
@@
For any $\epsilon,\delta\in(0,1/2)$ and $B>2$, a sparse subspace embedding matrix $\mathbf{\Pi}$ with $s=O(Bd\log(d/\delta)/\epsilon^{2})$ and $\gamma=O(\log_{B}(d/\delta)/\epsilon)$ satisfies
${\mathbb{E}}_{\mathbf{\Pi}\sim\mathcal{D}}\|(\mathbf{\Pi}\mathbf{U})^{\top}(\mathbf{\Pi}\mathbf{U})-\mathbf{I}\|^{\log(d/\delta)}<\epsilon^{\log(d/\delta)}\cdot\delta.$
@@
    approximate spectral norm matrix multiplication w.r.t. $\mathbf{U}_{k},\mathbf{A}_{\bar{k}}\coloneqq\mathbf{A}-\mathbf{A}_{k}$, that is:
@@
    $\mathbf{S}$ is a $1/2$-subspace embedding for the column space of $\mathbf{U}_{k}$
then by denoting $\tilde{\mathbf{A}}_{k}$ as the rank-$k$ approximation of $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{P}_{\mathbf{S}}$, we have
@@ 5 Three-Level MSP for Solving General Linear Systems
Since $\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}\in\mathcal{S}_{n}^{++}$, one straightforward idea is to construct the Nyström preconditioner similar as before. Given sketching matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$, define $\tilde{\mathbf{A}}\coloneqq\mathbf{A}\mathbf{S}^{\top},\mathbf{C}\coloneqq\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}=\mathbf{A}^{\top}\tilde{\mathbf{A}}$ and $\mathbf{W}\coloneqq\mathbf{S}\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}=\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}$, then $\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$ is the classical Nyström preconditioner of matrix $\mathbf{A}^{\top}\mathbf{A}$. Let $\{\sigma_{i}\}_{i=1}^{n}$ be the singular values of $\mathbf{A}$ in decreasing order. By using Lemma 4.2, similar guarantee holds if we set $\mathbf{M}\coloneqq\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}\coloneqq\lambda+\frac{2}{l}\sum_{i>l}\sigma_{i}^{2}$, and we have the following holds:
@@
However, different from the positive definite case where we can compute $\mathbf{C}$ and $\mathbf{W}$ explicitly, in this case constructing $\mathbf{C}$ and $\mathbf{W}$ takes $\tilde{O}(s\cdot\mathrm{nnz}(\mathbf{A}))$ time which is unaffordable. Instead, suppose we have the guarantee $\tilde{\mathbf{A}}\tilde{\mathbf{A}}^{\top}+\tilde{\lambda}\mathbf{I}\approx_{2}\mathbf{A}\mathbf{A}^{\top}+\tilde{\lambda}\mathbf{I}$ (which is proved in Lemma 5.1), then we have
@@
With this observation, we do not need to compute $\mathbf{C}$, and we can use $\mathbf{M}_{2}\coloneqq\mathbf{W}^{2}+\tilde{\lambda}\mathbf{W}$ as the Level-2 preconditioner for solving the linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{r}$. However we still cannot afford to compute $\mathbf{W}$, not to mention $\mathbf{M}_{2}$. Notice that
@@
since we only need to have access to $\mathbf{M}_{2}^{-1}\mathbf{r}$ in Level-2 preconditioned Lanczos, it suffices to compute $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})^{-1}\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\lambda\mathbf{I})^{-1}\mathbf{r}$. To solve this, we introduce the third level, where we again use preconditioned Lanczos to solve two positive definite linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$. To construct the Level-3 preconditioners, notice that $\tilde{\mathbf{A}}$ is an $m\times s$ tall matrix, and we can thus construct an $\tilde{O}(s)\times s$ sketch of it.
Let $\mathbf{\Phi}\in\mathbb{R}^{\phi\times m}$ be the sparse embedding matrix in Lemma 4.4 with $\phi=\tilde{O}(s)$. We first compute $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}\in\mathbb{R}^{\phi\times s}$, then construct the preconditioner $\mathbf{M}_{3a}^{-1}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}^{-1}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I}$. Lemma 4.4 guarantees that both preconditioned linear systems have a constant condition number, thus solving them using preconditioned Lanczos only takes $O(\log 1/\epsilon_{2})$ steps for the targeted accurancy $\epsilon_{2}$. Moreover, since both $\mathbf{M}_{3a}$ and $\mathbf{M}_{3b}$ are $s\times s$ positive definite matrices, we can afford to compute their inverse exactly, and the errors caused by the Level-3 preconditioners are 00. By going back from Level-3 $\to$ Level-2 $\to$ Level-1 and analyzing the error bounds carefully, we can finally arrive at the convergence result on the original linear system.
Our algorithms are as follows. Algorithm 3 is the main algorithm for solving the linear system $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{c}$, in which we run preconditioned Lanczos with Level-1 preconditioner $\mathbf{M}$ defined by function $\mathsf{SolverM1}$. Algorithm 4 defines function $\mathsf{SolverM1}$ by solving linear system $\mathbf{M}\mathbf{w}=\mathbf{r}$ with preconditioned Lanczos, where the Level-2 preconditioner $\mathbf{M}_{2}$ is defined by function $\mathsf{SolverM2}$. Lastly, Algorithm 5 defines function $\mathsf{SolveM2}$ by solving two linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$ with preconditioned Lanczos respectively, whereas the Level-3 preconditioners $\mathbf{M}_{3a}$ and $\mathbf{M}_{3b}$ are pre-computed (see line 5 of Algorithm 3). As a comparison of the quality of the preconditioners we use in different levels, we have:
@@
1:Input: matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, regularize term $\lambda$ and $\lambda_{0}$, vector $\mathbf{c}\in\mathbb{R}^{n}$, sparse sketch size $s$, # of non-zeros $\gamma$, # of iterations $t_{\max}$;
@@
3:Compute $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{S}^{\top}$ and $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}$; $\triangleright$ Lemma 4.4.
4:Compute $\mathbf{M}_{3a}=\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}=\mathbf{M}_{3a}+\tilde{\lambda}\mathbf{I}$
5:Compute $\mathbf{M}_{3a}^{-1}$ and $\mathbf{M}_{3b}^{-1}$;
6:Solve $\tilde{\mathbf{x}}$ by preconditioned Lanczos with $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I},\mathbf{c},\mathsf{SolveM1},t_{\max})$; $\triangleright$ Alg. 6.
7:return $\tilde{\mathbf{x}}$; $\triangleright$ Solves $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{c}$.
@@
4: Compute $\hat{\mathbf{z}}=\frac{1}{\lambda}(\hat{\mathbf{u}}-\hat{\mathbf{v}})$;
@@
As the innermost level (Algorithm 5), we use preconditioned Lanczos to solve two positive definite linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$, with preconditioners $\mathbf{M}_{3a}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I}$ respectively, where $\hat{\mathbf{A}}\coloneqq\mathbf{\Phi}\tilde{\mathbf{A}}\in\mathbb{R}^{\phi\times s}$ and $\mathbf{\Phi}$ is the sparse embedding matrix defined by Lemma 4.4 with $\phi=O(s+\log(1/\delta))$. With probability $31-\delta/$ 3 we have $\kappa_{\mathbf{M}_{3a}}=\kappa_{\mathbf{M}_{3b}}=O(1)$. Since in Algorithm 3 we pre-compute $\mathbf{M}_{3a}^{-1}$ and $\mathbf{M}_{3b}^{-1}$ exactly, the error terms for these two preconditioners are 00. By using Lemma 4.1, we need $O(\sqrt{\kappa_{\mathbf{M}_{3a}}}\log(\kappa_{\mathbf{M}_{3a}}/\epsilon_{2,1}))$ and $O(\sqrt{\kappa_{\mathbf{M}_{3b}}}\log(\kappa_{\mathbf{M}_{3b}}/\epsilon_{2,2}))$ iterations respectively to guarantee
@@
Finally we consider the total cost of Algorithm 3 together with Algorithm 4 and 5. We first pre-compute $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{S}^{\top}$ which takes $O(\mathrm{nnz}(\mathbf{A})\log(l/\delta))$, and also compute $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}$ which takes $O(\mathrm{nnz}(\tilde{\mathbf{A}})\log(s/\delta)+s^{2}\log^{4}(s/\delta))$ according to Lemma 4.4. With these, we can compute the Level-3 preconditioners $\mathbf{M}_{3a}^{-1}=(\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}})^{-1}$ and $\mathbf{M}_{3b}^{-1}=(\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I})^{-1}$ in time $O(s^{\omega})$. To conclude, line 2-5 of Algorithm 3 takes in total:

### v2

@@ 4 Two-Level MSP for Positive Definite Linear Systems
and for the other side, using that $\|\hat{\mathbf{A}}_{\mathrm{nys}}-\mathbf{A}\|\leq\lambda_{0}$ we have
@@
For the second level we use preconditioned Lanczos to solve the $s\times s$ linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{C}^{\top}\mathbf{r}$. As for the preconditioner we choose $\mathbf{M}_{2}=\tilde{\mathbf{C}}^{\top}\tilde{\mathbf{C}}+\tilde{\lambda}\mathbf{W}=\mathbf{C}^{\top}\mathbf{\Phi}^{\top}\mathbf{\Phi}\mathbf{C}+\tilde{\lambda}\mathbf{W}$, where $\mathbf{\Phi}\in\mathbb{R}^{\phi\times n}$ is a sparse sketching matrix constructed using standard subspace embedding techniques [CDDR24] (see Lemma 2.2) with choice $\phi=O(s+\log 1/\delta)$. By Lemma 2.2, with probability $21-\delta/$ 2 we have $\kappa_{\mathbf{M}_{2}}\coloneqq\kappa(\mathbf{M}_{2}^{-1/2}(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{M}_{2}^{-1/2})=O(1)$. Since in Algorithm 1 we pre-compute $\mathbf{M}_{2}^{-1}$ exactly, the error caused by the preconditioner is 00. By using Theorem 6.1 we know that after $O(\sqrt{\kappa_{\mathbf{M}_{2}}}\log(\kappa_{\mathbf{M}_{2}}/\epsilon_{1}))=O(\log 1/\epsilon_{1})=O(\log(\kappa/\epsilon_{0}))$ iterations, preconditioned Lanczos will output $\hat{\mathbf{y}}$ that satisfies
@@ 4.2 Coarse Nyström Preconditioner Based on Sparse Embedding
In this section we give a proof of Lemma 4.2, which measures the quality of Nyström preconditioner based on a sparse embedding matrix. Notice that, compared to the most commonly used guarantee that the preconditioned linear system has a constant condition number, we are using a “coarse” Nyström preconditioner which benefits us by reducing the cost of constructing it, while slightly sacrificing its quality and allowing the condition number after preconditioning to be “not so small”.
To characterize the properties of sparse embedding matrix, we first introduce the following notion of oblivious subspace embedding moments property, which can be seen as a generalization of the “JL-moment property” introduced by [KN14].
@@
[Coh16] showed that a sparse embedding matrix (Definition 2.1) with sketch size $s=O(d\log(d/\delta)/\epsilon^{2})$ and $\gamma=O(\log(d/\delta)/\epsilon)$ non-zeros per column satisfies $(\epsilon,\delta,d,\log(d/\delta))$-OSE moments property, as stated in the following lemma.
@@
For any $\epsilon,\delta\in(0,1/2)$ and $B>2$, a sparse subspace embedding matrix $\mathbf{S}$ with $s=O(Bd\log(d/\delta)/\epsilon^{2})$ and $\gamma=O(\log_{B}(d/\delta)/\epsilon)$ satisfies
${\mathbb{E}}\,\|(\mathbf{S}\mathbf{U})^{\top}(\mathbf{S}\mathbf{U})-\mathbf{I}\|^{\log(d/\delta)}<\epsilon^{\log(d/\delta)}\cdot\delta.$
@@
    approximate spectral norm matrix multiplication w.r.t. $\mathbf{U}_{k}$ and $\mathbf{A}_{\bar{k}}\coloneqq\mathbf{A}-\mathbf{A}_{k}$, that is:
@@
    $\mathbf{S}$ is a $1/2$-subspace embedding for the column space of $\mathbf{U}_{k}$,
then, defining $\mathbf{P}_{\mathbf{S}}$ as the orthogonal projection onto the row-space of $\mathbf{S}\mathbf{A}$, and denoting $\tilde{\mathbf{A}}_{k}$ as the best rank-$k$ approximation of $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{P}_{\mathbf{S}}$, we have:
@@ 5 Three-Level MSP for Solving General Linear Systems
Since $\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I}\in\mathcal{S}_{n}^{++}$, one natural idea is to construct the Nyström preconditioner similar as before. Given sketching matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$, define $\tilde{\mathbf{A}}\coloneqq\mathbf{A}\mathbf{S}^{\top},\mathbf{C}\coloneqq\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}=\mathbf{A}^{\top}\tilde{\mathbf{A}}$ and $\mathbf{W}\coloneqq\mathbf{S}\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}=\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}$. Then, $\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$ is the classical Nyström preconditioner of matrix $\mathbf{A}^{\top}\mathbf{A}$. Let $\{\sigma_{i}\}_{i=1}^{n}$ be the singular values of $\mathbf{A}$ in decreasing order. By using Lemma 4.2, we can again show that the matrix $\mathbf{M}\coloneqq\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}+\tilde{\lambda}\mathbf{I}$ where $\tilde{\lambda}\coloneqq\lambda+\frac{2}{l}\sum_{i>l}\sigma_{i}^{2}$, is a useful preconditioner for this problem, and we also have the inversion formula:
@@
However, different from the positive definite case where we can compute $\mathbf{C}$ and $\mathbf{W}$ explicitly, in this case constructing $\mathbf{C}$ and $\mathbf{W}$ takes $\tilde{O}(s\cdot\mathrm{nnz}(\mathbf{A}))$ time which is much too expensive. To avoid computing $\mathbf{C}$, we first observe that $\tilde{\mathbf{A}}\tilde{\mathbf{A}}^{\top}+\tilde{\lambda}\mathbf{I}\approx_{2}\mathbf{A}\mathbf{A}^{\top}+\tilde{\lambda}\mathbf{I}$ (which is proved in Lemma 5.1), concluding that
@@
With this observation, we do not need to compute $\mathbf{C}$, and we can use $\mathbf{M}_{2}\coloneqq\mathbf{W}^{2}+\tilde{\lambda}\mathbf{W}$ as the Level-2 preconditioner for solving the linear system $(\mathbf{C}^{\top}\mathbf{C}+\tilde{\lambda}\mathbf{W})\mathbf{y}=\mathbf{r}$. However we still cannot afford to compute $\mathbf{W}$, not to mention $\mathbf{M}_{2}$. To address this, notice that
@@
$=$ $~{}\frac{1}{\tilde{\lambda}}\left((\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})^{-1}-(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})^{-1}\right).$
Since we only need to have access to $\mathbf{M}_{2}^{-1}\mathbf{r}$ in Level-2 preconditioned Lanczos, it suffices to compute $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})^{-1}\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\lambda\mathbf{I})^{-1}\mathbf{r}$. To solve this, we introduce the third level, where we again use preconditioned Lanczos to solve two positive definite linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$. To construct the Level-3 preconditioners, notice that $\tilde{\mathbf{A}}$ is an $m\times s$ tall matrix, and we can thus construct an ${O}(s)\times s$ sketch of it.
Let $\mathbf{\Phi}\in\mathbb{R}^{\phi\times m}$ be the oblivious embedding matrix in Lemma 2.2 with $\phi={O}(s)$. We first compute $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}\in\mathbb{R}^{\phi\times s}$, then construct the preconditioner $\mathbf{M}_{3a}^{-1}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}^{-1}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I}$. Lemma 2.2 guarantees that both preconditioned linear systems have a constant condition number, thus solving them using preconditioned Lanczos only takes $O(\log 1/\epsilon_{2})$ steps for the targeted accurancy $\epsilon_{2}$. Moreover, since both $\mathbf{M}_{3a}$ and $\mathbf{M}_{3b}$ are $s\times s$ positive definite matrices, we can afford to compute their inverse exactly, and the errors caused by the Level-3 preconditioners are 00. By going back from Level-3 $\to$ Level-2 $\to$ Level-1 and analyzing the error bounds carefully, we can finally arrive at the convergence result on the original linear system.
Our algorithms are as follows. Algorithm 3 is the main algorithm for solving the linear system $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{c}$, in which we run preconditioned Lanczos with Level-1 preconditioner $\mathbf{M}$ defined by function $\mathsf{SolverM1}$. Algorithm 4 defines function $\mathsf{SolverM1}$ by solving linear system $\mathbf{M}\mathbf{w}=\mathbf{r}$ with preconditioned Lanczos, where the Level-2 preconditioner $\mathbf{M}_{2}$ is defined by function $\mathsf{SolverM2}$. Lastly, Algorithm 5 defines function $\mathsf{SolveM2}$ by solving two linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$ with preconditioned Lanczos respectively, whereas the Level-3 preconditioners $\mathbf{M}_{3a}$ and $\mathbf{M}_{3b}$ are pre-computed (see line 6 of Algorithm 3). As a comparison of the quality of the preconditioners we use in different levels, we have:
@@
1:Input: matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, regularize term $\lambda$ and $\lambda_{0}$, vector $\mathbf{c}\in\mathbb{R}^{n}$, sparse sketch size $s$, # of non-zeros $\gamma$, level-2 sketch size $\phi$, # of iterations $t_{\max}$;
@@
3:Construct sparse embedding matrix $\mathbf{S}\in\mathbb{R}^{s\times n}$ with $\gamma$ non-zeros per column; $\triangleright$ Def. 2.1.
4:Compute $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{S}^{\top}$ and $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}\in\mathbb{R}^{\phi\times s}$; $\triangleright$ Lemma 2.2.
5:Compute $\mathbf{M}_{3a}=\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}=\mathbf{M}_{3a}+\tilde{\lambda}\mathbf{I}$
6:Compute $\mathbf{M}_{3a}^{-1}$ and $\mathbf{M}_{3b}^{-1}$;
7:Solve $\tilde{\mathbf{x}}$ by preconditioned Lanczos with $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I},\mathbf{c},\mathsf{SolveM1},t_{\max})$; $\triangleright$ Alg. 6.
8:return $\tilde{\mathbf{x}}$; $\triangleright$ Solves $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{c}$.
@@
4: Compute $\hat{\mathbf{z}}=\frac{1}{\tilde{\lambda}}(\hat{\mathbf{u}}-\hat{\mathbf{v}})$;
@@
As the innermost level (Algorithm 5), we use preconditioned Lanczos to solve two positive definite linear systems $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}})\mathbf{u}=\mathbf{r}$ and $(\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}+\tilde{\lambda}\mathbf{I})\mathbf{v}=\mathbf{r}$, with preconditioners $\mathbf{M}_{3a}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}$ and $\mathbf{M}_{3b}\coloneqq\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I}$ respectively, where $\hat{\mathbf{A}}\coloneqq\mathbf{\Phi}\tilde{\mathbf{A}}\in\mathbb{R}^{\phi\times s}$ and $\mathbf{\Phi}$ is the oblivious embedding matrix defined by Lemma 2.2 with $\phi=O(s+\log(1/\delta))$. With probability $31-\delta/$ 3 we have $\kappa_{\mathbf{M}_{3a}}=\kappa_{\mathbf{M}_{3b}}=O(1)$. Since in Algorithm 3 we pre-compute $\mathbf{M}_{3a}^{-1}$ and $\mathbf{M}_{3b}^{-1}$ exactly, the error terms for these two preconditioners are 00. By using Lemma 4.1, we need $O(\sqrt{\kappa_{\mathbf{M}_{3a}}}\log(\kappa_{\mathbf{M}_{3a}}/\epsilon_{2,1}))$ and $O(\sqrt{\kappa_{\mathbf{M}_{3b}}}\log(\kappa_{\mathbf{M}_{3b}}/\epsilon_{2,2}))$ iterations respectively to guarantee
@@
Finally we consider the total cost of Algorithm 3 together with Algorithm 4 and 5. We first pre-compute $\tilde{\mathbf{A}}=\mathbf{A}\mathbf{S}^{\top}$ which takes $O(\mathrm{nnz}(\mathbf{A})\log(l/\delta))$, and also compute $\hat{\mathbf{A}}=\mathbf{\Phi}\tilde{\mathbf{A}}$ which takes $O(\mathrm{nnz}(\tilde{\mathbf{A}})\log(s/\delta)+s^{2}\log^{4}(s/\delta))$ according to Lemma 2.2. With these, we can compute the Level-3 preconditioners $\mathbf{M}_{3a}^{-1}=(\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}})^{-1}$ and $\mathbf{M}_{3b}^{-1}=(\hat{\mathbf{A}}^{\top}\hat{\mathbf{A}}+\tilde{\lambda}\mathbf{I})^{-1}$ in time $O(s^{\omega})$. To conclude, line 2-6 of Algorithm 3 takes in total: